# RxnCaption: Reformulating Reaction Diagram Parsing as Visual Prompt Guided Captioning

### v1

@@ Abstract
Large-scale chemical reaction datasets are crucial for AI research in chemistry. However, existing chemical reaction data often exist as images within papers, making them not machine-readable and unusable for training machine learning models. In response to this challenge, we propose the RxnCaption framework for the task of chemical Reaction Diagram Parsing (RxnDP). Our framework reformulates the traditional coordinate prediction driven parsing process into an image captioning problem, which Large Vision-Language Models (LVLMs) handle naturally. We introduce a strategy termed “BBox and Index as Visual Prompt” (BIVP), which uses our state-of-the-art molecular detector, MolYOLO, to pre-draw molecular bounding boxes and indices directly onto the input image. This turns the downstream parsing into a natural-language description problem. Extensive experiments show that the BIVP strategy significantly improves structural extraction quality while simplifying model design. We further construct the RxnCaption-11k dataset, an order of magnitude larger than prior real-world literature benchmarks, with a balanced test subset across four layout archetypes. Experiments demonstrate that RxnCaption-VL achieves state-of-the-art performance on multiple metrics. We believe our method, dataset, and models will advance structured information extraction from chemical literature and catalyze broader AI applications in chemistry. We will release data, models, and code on GitHub.
@@ 1 Introduction
Machine learning excels in predicting chemical reaction outcomes [31], optimizing conditions [7], and planning retrosynthetic routes [46]. However, progress is limited by the scarcity of large, diverse datasets [28, 46, 8]. Public datasets like USPTO-50k [30] lack the size and diversity of commercial databases such as SciFinder [10] and Reaxys [17]. Additionally, much valuable data is trapped in unstructured image formats in scientific literature, hindering machine learning training. Automated deep learning approaches offer a promising solution, with significant progress in molecular structure image recognition [26] and text-based reaction extraction [45]. The use of Large Visual Language Models (LVLMs) [37] has further accelerated advancements in this area.
@@
Recent studies [39, 25, 5] have applied deep learning to RxnDP. RxnScribe [25] uses the Pix2Seq [4] model for simultaneous molecule detection and reaction extraction via a sequence generation approach, termed as the “Bbox and Role in One Step” (BROS) strategy. RxnIm [5] was the first to apply the LVLM to RxnDP, also using the BROS strategy and significantly increasing training data through data synthesis. However, its performance improvement over RxnScribe is modest, and it struggles with out-of-distribution (OOD) generalization. Thus, the application of LVLMs in this field has not yet led to the breakthroughs seen in other domains.
To achieve significant progress in RxnDP with LVLMs, improvements are needed in two areas, addressing RxnIm’s limitations. Firstly, the BROS strategy may not suit LVLMs, as coordinate prediction falls outside their core competencies [19, 14]. Instead, prediction strategies should align with LVLMs’ inherent strengths. We propose the RxnCaption framework, which allows LVLMs to parse reactions by describing images, leveraging their natural language generation capabilities. Using a visual prompt approach [40], we employ MolYOLO, a specialized molecular detection model we developed, to pre-draw molecular bounding boxes and indice on the image, creating a pre-annotated image. The LVLM then parses the reaction by describing the image, referencing molecular indices, and using OCR for text recognition. This “BBox and Index as Visual Prompt” (BIVP) strategy enables parsing entirely in natural language, maximizing their potential. Secondly, from a data perspective, existing RxnDP datasets are limited: the RxnScribe-train set contains only a few samples from real papers, while the RxnIm-train set, though larger, is synthetic and differs from real images, hindering model generalization. To address this, we created the RxnCaption-11k dataset, a large, high-quality, diverse RxnDP dataset.
Using the RxnCaption framework and RxnCaption-11k dataset, we developed the RxnCaption-VL model, achieving state-of-art performance on both RxnScribe-test and RxnCaption-11k-test datasets. We hope that our methods, data, and models will advance RxnDP technology and support the development of AI in chemistry.
@@
2. We created RxnCaption-11k, the largest dataset for chemical reaction diagram parsing from real papers, which is an order of magnitude larger than RxnScribe dataset.
@@ 2.2 Reflection on the BROS Strategy
Initially, we evaluated the RxnDP performance of leading LVLMs using the approach from RxnScribe [25] and RxnIm [5]. These models were tasked with predicting both the bounding boxes (bboxes) of molecular components and their roles simultaneously, in a method we term ”Bbox and Role in One Step” (BROS).
We tested models such as GPT-4o [13], Gemini-2.5-Pro [33], and Qwen-VL-max [1] in a zero-shot scenario on a subset of the RxnScribe-test set, labeled RxnScribe-test-slct. [^1] This subset includes only samples where all molecules participate in the reaction and have complete bbox annotations, excluding interference from free molecules. Details on the prompt templates and evaluation metrics are in Appendix B and § 4.1. As shown in Table 1, despite their strong performance in general tasks, these models perform poorly on RxnDP using the BROS strategy.
This is unexpected, given the extensive interleaved text and image data from chemical literature and web sources in LVLM pre-training datasets [20]. To delve deeper, we devised a visual question answering (VQA) task to assess understanding of reaction diagrams, with questions like “How many chemical reactions are depicted?” and “Does the diagram include cyclic structures?” (details in Appendix B). As shown in Table 1, mainstream LVLMs exhibit a good understanding of reaction diagrams, with Gemini-2.5 Pro achieving an accuracy of 75.9%.
Comparative analysis indicates that the core issue is the LVLMs’ need to predict molecular bboxes. The BROS strategy lies beyond LVLMs’ inherent strengths [19, 14]. RxnIm encountered similar challenges with BROS, addressing them through multi-stage training, initially using synthetic data to enhance molecular bbox prediction. However, this approach is complex and costly, with limited performance gains and a lack of generalization capability (see § 4.2).
@@ 2.3 Our Insights and Preliminary Verification
The VQA results confirm that mainstream VLMs can effectively understand chemical reaction diagrams, indicating they possess relevant domain knowledge and capabilities. This leads to a key insight: instead of forcing LVLMs to learn new skills during fine-tuning, we should develop prediction strategies that leverage their inherent capabilities. Inspired by research on visual prompts [40], we introduce the BBox and Index as Visual Prompt (BIVP) strategy:
@@ 3 Method
Figure 3: Construction pipeline of RxnCaption-11k dataset.
@@ 3.1.2 Training and Performance
We based our detector on the YOLOv10 [36] architecture, using the MolDet-33k-train set and the YoDe dataset [47] (1,738 images, 15,887 bboxes). We evaluated MolYOLO, YoDe, MolDetect [^2] https://github.com/Ozymandias314/MolDetect on MolDet-33k-test. As shown in Table 2, MolYOLO achieved state-of-the-art performance, with particularly high accuracy and recall, demonstrating its strong detection capabilities.
(a) Layout distribution of RxnScribe-test and RxnCaption-11k-test.
(b) t-SNE visualization of the image embeddings for RxnScribe, RxnIm, and RxnCaption-11k.
Figure 4: (a) t-SNE visualization of dataset embeddings; (b) Layout distribution comparison between RxnScribe and RxnCaption-11k.
3.2 RxnCaption-11k Dataset
Existing RxnDP datasets have notable limitations: RxnScribe dataset, derived from real papers, includes only 1378 samples, while RxnIm dataset, though large, is artificially synthesized, limiting diversity and model generalization (see § 4.2). To overcome these issues, we developed the RxnCaption-11k dataset—a large, high-quality, and diverse RxnDP dataset.
@@ 3.2.1 Construction
RxnCaption-11k uses the same chemical paper sources and Fig/Tab Scope images as MolDet-33k. Its construction involves four steps (Figure 3): (1) Molecular Structure Annotation: Annotate molecular structure bboxes, completed during MolDet-33k construction. (2) Reaction Region Annotation: Use irregular polygons to mark each chemical reaction area. (3) Component Role Annotation: Annotate the chemical role of each reaction component, with text components following RxnScribe guidelines. (4) Text Content Extraction: Training Set: Use Gemini-2.5 Pro’s OCR for automated text extraction (details in Appendix C). Validation Set: Use manual annotation for accuracy.
@@ 3.2.2 Statistics and Comparison
Table 3: Statistics of the RxnDP datasets. ‘ours’ means RxnCaption-11k.
Following MolDet-33k’s division strategy, samples from papers published by June 2024 were used for training, and those from July 2024 for validation. Chemical reaction diagrams are categorized into Single-line, Multi-line, Tree, and Cyclic. Figure 4(a) shows that the RxnScribe-test set is imbalanced, with the Single-line category dominating at 50%, while Tree/Cyclic categories are only 23.2%/7.2%. To address this, RxnCaption-11k-test set was carefully selected to ensure category diversity, with 100 Fig/Tab images from each category. The final dataset includes: Training Set: 10,112 Fig/Tab images, 24,716 chemical reactions. Validation Set: 400 Fig/Tab images, 2,829 chemical reactions.
To evaluate the distribution of RxnCaption-11k, RxnScribe, and RxnIm datasets, we used the CLIP visual encoder [27] for feature extraction and visualized their distribution with t-SNE [22] (see Figure 4(b)). The results indicate: (1) Real Data Diversity RxnCaption-11k and RxnScribe datasets, both from real papers, show significant feature overlap. However, RxnCaption-11k’s broader coverage confirms its advantage in quantity and diversity. (2) Synthetic Data Domain Shift RxnIm dataset’s distribution shows little overlap with real datasets, highlighting a domain gap. This explains RxnIm’s performance bottleneck: despite large data size, its performance improves litte on the RxnScribe-test set, and even performs worse on RxnCaption-11k compared to RxnScribe (see § 4.2).
@@ 3.3 RxnCaption-VL
We expanded the RxnScribe-train set from 1,240 to 4,368 images and combined it with the RxnCaption -11k-train set of 10,112 images, creating a training set of 17,442, images. We fine-tuned the Qwen2.5-VL-7B model, resulting our RxnCaption-VL model.
@@ 4 Experiments
Model Strategy RxnScribe-test RxnCaption-11k-test
@@
RxnCaption-VL BIVP 76.0 75.0 75.5 88.5 87.8 88.2 57.3 55.6 55.5 68.7 66.6 67.6
BROS 74.1 69.9 71.9 80.8 77.2 78.9 54.3 52.0 53.2 64.6 61.5 63.0
RxnScribe_w/11k BROS 76.3 69.1 72.6 85.8 78.0 81.7 60.7 35.6 44.8 71.8 42.6 53.5
RxnScribe_offical BROS 76.2 71.9 74.0 86.1 81.6 83.8 47.4 27.6 34.9 62.1 36.4 45.9
RxnIM BROS 75.4 71.2 73.2 79.2 74.7 76.9 48.8 30.3 37.4 52.9 32.8 40.5
@@
Close-source Model
@@
Table 4: Model performance. ‘RxnScribe_offical’ means the offically released model from [25], ‘RxnScribe_w/11k’ means we retrain the RxnScribe with the same training data as RxnCaption-VL using BROS strategy.
@@ 4.1.2 Baselines
We select two types of baseline models: the first type uses current mainstream general VLMs, including GPT4o, Gemini-2.5-Pro (all following metrics were tested in October 2025), and Qwen2.5-VL and Intern-VL3; the second type is the specialized models for RxnDP task, primarily RxnScribe and RxnIm.
During testing, we compare the BROS and BIVP strategies for the general VLM models, while the specialized models continue to utilize their native BROS strategy.
@@ 4.1.3 Metric
We follows the evaluation framework of RxnScribe [25], constructing an assessment system at the chemical reaction granularity. Given the predicted reaction set $\mathcal{R}^{\text{pred}}=\{R_{i}^{\text{pred}}\}$ and the ground truth reaction set $\mathcal{R}^{\text{gt}}=\{R_{i}^{\text{gt}}\}$, performance metrics are calculated through reaction instance matching: a match between instances in the predicted and true sets is recorded as a true positive (TP), unmatched predicted reactions are false positives (FP), and unmatched true reactions are false negatives (FN).
To evaluate performance, we utilize two complementary matching strategies: SoftMatch and HybridMatch. SoftMatch follows the RxnScribe benchmark protocol by excluding text components from evaluation and integrating condition components into reactant components. A reaction is classified as a TP only if all molecular bounding boxes within the reactants and products have an Intersection over Union (IoU) greater than 0.5. This strategy is applicable to both BROS and BIVP methodologies. Conversely, HybridMatch is tailored to accommodate variations in output representations based on the specific strategy used. For BROS, it applies a rigorous HardMatch criterion consistent with the RxnScribe standard, requiring an IoU greater than 0.5 for all components, including reactants, products, and conditions. For BIVP, molecular components must have exact bounding box matches with an IoU exceeding 0.5, while text components are verified through content-based methods. A match of text component is confirmed if the edit distance ratio to length between the ground truth and predicted text strings is no more than 0.2. Finally, based on the TP/FP/FN statistics, the precision, recall, and F1 score under both SoftMatch and HybridMatch are calculated.

### v2

@@ Abstract
Large-scale chemical reaction datasets are crucial for AI research in chemistry. However, existing chemical reaction data often exist as images within papers, making them not machine-readable and unusable for training machine learning models. In response to this challenge, we propose the RxnCaption framework for the task of chemical Reaction Diagram Parsing (RxnDP). Our framework reformulates the traditional coordinate prediction driven parsing process into an image captioning problem, which Large Vision-Language Models (LVLMs) handle naturally. We introduce a strategy termed “BBox and Index as Visual Prompt” (BIVP), which uses our state-of-the-art molecular detector, MolYOLO, to pre-draw molecular bounding boxes and indices directly onto the input image. This turns the downstream parsing into a natural-language description problem. Extensive experiments show that the BIVP strategy significantly improves structural extraction quality while simplifying model design. We further construct the RxnCaption-15k dataset, an order of magnitude larger than prior real-world literature benchmarks, with a balanced test subset across four layout archetypes. Experiments demonstrate that RxnCaption-VL achieves state-of-the-art performance on multiple metrics. We believe our method, dataset, and models will advance structured information extraction from chemical literature and catalyze broader AI applications in chemistry. We will release data, models, and code on GitHub.
@@ 1 Introduction
Machine learning excels in predicting chemical reaction outcomes [schwaller2019molecular], optimizing conditions [de2019synthetic], and planning retrosynthetic routes [zhong2024recent]. However, progress is limited by the scarcity of large, diverse datasets [ramos2025review, zhong2024recent, ding2025survey]. Public datasets like USPTO-50k [schneider2016s] lack the size and diversity of commercial databases such as SciFinder [gabrielson2018scifinder] and Reaxys [lawson2014making]. Additionally, much valuable data is trapped in unstructured image formats in scientific literature, hindering machine learning training. Automated deep learning approaches offer a promising solution, with significant progress in molecular structure image recognition [qian2023molscribe] and text-based reaction extraction [zhong2023reaction]. The use of Large Visual Language Models (LVLMs) [wang2025gtr] has accelerated advancements in this area.
@@
Recent studies [wilary2023reactiondataextractor, qian2023rxnscribe, RxnIm] have applied deep learning to RxnDP. RxnScribe [qian2023rxnscribe] uses the Pix2Seq [chen2021pix2seq] model for simultaneous molecule detection and reaction extraction via a sequence generation approach, termed as the “Bbox and Role in One Step” (BROS) strategy. RxnIM [RxnIm] was the first to apply the LVLM to RxnDP, also using the BROS strategy and significantly increasing training data through data synthesis. However, its performance improvement over RxnScribe is modest, and it struggles with out-of-distribution generalization. Thus, the application of LVLMs in this field has not yet led to the breakthroughs seen in other domains.
To achieve significant progress in RxnDP with LVLMs, improvements are needed in two areas, addressing RxnIM’s limitations. Firstly, the BROS strategy may not suit LVLMs, as coordinate prediction falls outside their core competencies [li2025lmm, kang2025your]. Instead, prediction strategies should align with LVLMs’ inherent strengths. We propose the RxnCaption framework, which allows LVLMs to parse reactions by describing images, leveraging their natural language generation capabilities. Using a visual prompt approach [wu2024visual], we employ MolYOLO, a specialized molecular detection model we developed, to pre-draw molecular bounding boxes and indice on the image, creating a pre-annotated image. The LVLM then parses the reaction by describing the image, referencing molecular indices, and using OCR for text recognition. This “BBox and Index as Visual Prompt” (BIVP) strategy enables parsing entirely in natural language, maximizing their potential. Secondly, from a data perspective, existing RxnDP datasets are limited: the RxnScribe-train set contains only a few samples from real papers, while the RxnIM-train set, though larger, is synthetic and differs from real images, hindering model generalization. To address this, we created the RxnCaption-15k dataset, a large, high-quality, diverse RxnDP dataset.
Using the RxnCaption framework and RxnCaption-15k dataset, we developed the RxnCaption-VL model, achieving state-of-the-art performance on both RxnScribe-test and RxnCaption-15k-test datasets. We hope that our methods, data, and models will advance RxnDP task and support the development of AI for chemistry.
@@
2. We created RxnCaption-15k, the largest dataset for chemical reaction diagram parsing from real papers, an order of magnitude larger than the RxnScribe dataset.
@@ 2.2 Reflection on the BROS Strategy
Initially, we evaluated the RxnDP performance of leading LVLMs using the approach from RxnScribe [qian2023rxnscribe] and RxnIm [RxnIm]. These models were tasked with predicting both the bounding boxes (bboxes) of molecular components and their roles simultaneously, in a method we term ”Bbox and Role in One Step” (BROS).
We tested models such as GPT-4o [hurst2024gpt], Gemini-2.5-Pro [team2023gemini], and Qwen-VL-max [bai2023qwentechnicalreport] in a zero-shot scenario on a subset of the RxnScribe-test set, labeled RxnScribe-test-slct. [^1] This subset includes only samples where all molecules participate in the reaction and have complete bbox annotations, excluding interference from free molecules. Details on the prompt templates and evaluation metrics are in Appendix B and § 4.1. As shown in Table 1, despite their strong performance in general tasks, these models perform poorly on RxnDP using the BROS strategy.
This is unexpected, given the extensive interleaved text and image data from chemical literature and web sources in LVLM pre-training datasets [li2024omnicorpus]. To delve deeper, we devised a visual question answering (VQA) task to assess understanding of reaction diagrams, with questions like “How many chemical reactions are depicted?” and “Does the diagram include cyclic structures?” (details in Appendix B). As shown in Table 1, mainstream LVLMs exhibit a good understanding of reaction diagrams, with Gemini-2.5 Pro achieving an accuracy of 75.9%.
Comparative analysis indicates that the core issue is the LVLMs’ need to predict molecular bboxes. The BROS strategy lies beyond LVLMs’ inherent strengths [li2025lmm, kang2025your]. RxnIm encountered similar challenges with BROS, addressing them through multi-stage training, initially using synthetic data to enhance molecular bbox prediction. However, this approach is complex and costly, with limited performance gains and a lack of generalization capability (see § 4.2).
@@ 2.3 Our Insights and Preliminary Verification
The VQA results confirm that mainstream VLMs can effectively understand chemical reaction diagrams, indicating they possess relevant domain knowledge and capabilities. This leads to a key insight: instead of forcing LVLMs to learn new skills during fine-tuning, we should develop prediction strategies that leverage their inherent capabilities. Inspired by research on visual prompts [wu2024visual], we introduce the BBox and Index as Visual Prompt (BIVP) strategy:
@@ 3 Method
Figure 3: Construction pipeline of RxnCaption-15k dataset.
@@ 3.1.2 Training and Performance
We based our detector on the YOLOv10 [wang2024yolov10] architecture, using the MolDet-33k-train set and the YoDe dataset [zhou2023yode] (1,738 images, 15,887 bboxes). We evaluated MolYOLO, YoDe, MolDetect [^2] https://github.com/Ozymandias314/MolDetect on MolDet-33k-test. As shown in Table 2, MolYOLO achieved state-of-the-art performance, with particularly high accuracy and recall, demonstrating its strong detection capabilities.
(a) Layout distribution of RxnScribe-test and RxnCaption-15k-test.
(b) t-SNE visualization of the image embeddings for RxnScribe, RxnIm, and RxnCaption-15k.
Figure 4: (a) t-SNE visualization of dataset embeddings; (b) Layout distribution comparison between RxnScribe and RxnCaption-15k.
3.2 RxnCaption-15k Dataset
Existing RxnDP datasets have notable limitations: RxnScribe dataset, derived from real papers, includes only 1378 samples, while RxnIM dataset, though large, is artificially synthesized, limiting diversity and model generalization (see § 4.2). To overcome these issues, we developed the RxnCaption-15k dataset—a large, high-quality, and diverse RxnDP dataset.
@@ 3.2.1 Construction
RxnCaption-15k uses the same chemical paper sources and Fig/Tab-level images as MolDet-33k. Its construction involves four steps (Figure 3): (1) Molecular Structure Annotation: Annotate molecular structure bboxes, completed during MolDet-33k construction. (2) Reaction Region Annotation: Use irregular polygons to mark each chemical reaction area. (3) Component Role Annotation: Annotate the chemical role of each reaction component, with text components following RxnScribe guidelines. (4) Text Content Extraction: For training Set, we use Gemini-2.5 Pro’s OCR for automated text extraction (details in Appendix C). For validation set, we use manual annotation for accuracy.
@@ 3.2.2 Statistics and Comparison
Table 3: Statistics of the RxnDP datasets. ‘ours’ means RxnCaption-15k.
Following MolDet-33k’s division strategy, samples from papers published by June 2024 were used for training, and those from July 2024 for validation. Chemical reaction diagrams are categorized into Single-line, Multi-line, Tree, and Cyclic. Figure 4(a) shows that the RxnScribe-test set is imbalanced, with the Single-line category dominating at 50%, while Tree/Cyclic categories are only 23.2%/7.2%. To address this, RxnCaption-15k-test set was carefully selected to ensure category diversity, with 100 Fig/Tab images from each category. The final dataset statistics are summarized in Table 3.
To evaluate the distribution of RxnCaption-15k, RxnScribe, and RxnIM datasets, we used the CLIP visual encoder [radford2021learning] for feature extraction and visualized their distribution with t-SNE [maaten2008visualizing] (see Figure 4(b)). The results indicate: (1) Real Data Diversity RxnCaption-15k and RxnScribe datasets, both from real papers, show significant feature overlap. However, RxnCaption-15k’s broader coverage confirms its advantage in quantity and diversity. (2) Synthetic Data Domain Shift RxnIM dataset’s distribution shows little overlap with real datasets, highlighting a domain gap. This explains RxnIM’s performance bottleneck: despite large data size, its performance improves litte on the RxnScribe-test set, and even performs worse on RxnCaption-15k compared to RxnScribe (see § 4.2).
@@ 3.3 RxnCaption-VL
We expanded the RxnScribe-train set from 1,240 to 3,720 images to balance the ratio of and combined it with the RxnCaption-15k-train set of 15,128 images. We additionally applied double augmentation to images containing reversible reactions, right-to-left reactions, and bottom-to-top reactions, resulting in a training set of 23,432 images. Using this, we fine-tuned the Qwen2.5-VL-7B, resulting our RxnCaption-VL model.
@@ 4 Experiments
Table 4: Model performance comparison on the RxnScribe-test and RxnCaption-15k-test datasets. “RxnScribe_official” refers to the model from [qian2023rxnscribe]. “RxnScribe_w/15k” is the RxnScribe model retrained with our dataset. Best scores are in bold, second best are underlined.
Model Strategy RxnScribe-test RxnCaption-15k-test
@@
RxnCaption-VL BIVP 71.6 72.7 72.2 85.3 87.1 86.2 60.3 59.3 59.8 71.3 69.4 70.4
BROS 69.6 68.9 69.2 76.2 76.2 76.2 57.0 57.5 57.2 66.4 67.4 66.9
RxnScribe_w/15k BROS 72.4 69.1 70.7 84.1 81.7 82.8 61.2 38.7 47.4 72.1 44.7 55.2
RxnScribe_official BROS 72.3 66.2 69.1 83.8 76.5 80.0 47.4 27.6 34.9 62.1 36.4 45.9
RxnIM BROS 71.0 70.1 70.5 79.2 74.7 76.9 48.8 30.3 37.4 52.9 32.8 40.5
@@
Closed-source Model
@@ 4.1.2 Baselines
We select two types of baseline models: the first type uses current mainstream general VLMs under zero-shot setting, including GPT4o, Gemini-2.5-Pro (all following metrics were tested in October 2025), and Qwen2.5-VL and Intern-VL3; the second type is the specialized models for RxnDP task, primarily RxnScribe and RxnIM.
During testing, we compare the BROS and BIVP strategies for the general VLMs, while the specialized models continue to utilize their native BROS strategy.
@@ 4.1.3 Metric
We follow the evaluation framework of RxnScribe [qian2023rxnscribe], constructing an assessment system at the chemical reaction granularity. Given the predicted reaction set $\mathcal{R}^{\text{pred}}=\{R_{i}^{\text{pred}}\}$ and the ground truth reaction set $\mathcal{R}^{\text{gt}}=\{R_{i}^{\text{gt}}\}$, performance metrics are calculated through reaction instance matching: a match between instances in the predicted and true sets is recorded as a true positive (TP), unmatched predicted reactions are false positives (FP), and unmatched true reactions are false negatives (FN).
To evaluate model performance, we employ two complementary strategies: SoftMatch and HybridMatch. SoftMatch follows the RxnScribe protocol by excluding text components and merging condition components into reactants. A reaction is counted as a TP only if all molecular bounding boxes in reactants and products achieve an IoU $\geq$ 0.5. This applies to both BROS and BIVP.
HybridMatch adapts to differences in output formats across methods. For BROS, it uses the strict HardMatch rule from RxnScribe, requiring all molecular components (reactants, products, and conditions) to reach IoU $\geq$ 0.5. If text components fail IoU matching, a textual matching strategy is applied (will be described as follows). For BIVP, molecular components must have exact bounding-box matches (IoU $\geq$ 0.5). Text within reactants and products must match exactly, while condition text is considered correct if its normalized edit distance $\leq$ 0.2. Overall, BIVP is evaluated more strictly, since BROS considers text matched if either IoU or text content matches.
Finally, based on the TP/FP/FN statistics, we compute the precision, recall, and F1 scores under both SoftMatch and HybridMatch.